OpenStudy (anonymous):

a projectile travels 1120 feet per second. how fast is the projectile traveling in miles per hour?

OpenStudy (anonymous):

\[\frac{ 1120 ft }{ \sec } \times \frac{ 1 mile }{ 5280 ft} \times \frac{ 3600 \sec}{ 1 hour }\]

OpenStudy (anonymous):

solving that gives, 763.636363 mi/hr